# A Persistence-Driven Edit Distance for Graphs with Abstract Weights

### v3

@@ B Proofs
$W_{1}(D,D^{\prime})=W_{1}(\sum_{x\in D}\{x\},\sum_{y\in D^{\prime}}\{y\})=\sup_{k\in\text{Lip}(\mathbb{R}^{2}_{\Delta})}(\sum_{x\in D}k(x)-\sum_{y\in D^{\prime}}k(y)).$
@@
  • replacing $D^{\prime}$ with the empty diagram and noting:
    $\sup_{k\in\text{Lip}(\mathbb{R}^{2}_{\Delta})}(\sum_{x\in D}k(x))=\sum_{x\in D}\min_{y\in\Delta}\parallel x-y\parallel_{1};$
  • observing that when we have $W_{1}(D+D^{\prime\prime},D^{\prime}+D^{\prime\prime})$ the contributions by the elements in $D^{\prime\prime}$ when computing:
    $\sum_{x\in D+D^{\prime\prime}}k(x)-\sum_{y\in D^{\prime}+D^{\prime\prime}}k(y)$
@@
Proof of 3.
We prove $s+s^{\prime}\in\mathcal{S}(G)$ and that it is well defined on the equivalence classes of $\mathcal{S}(G)$.
Because of the intersection condition then $s+s^{\prime}$ is injective. By construction it is also continuous. Moreover it is automatically an homeomorphism onto its image in a small enough neighbourhood of any point except for $1/2$. Let $(a,b)$ be an open neighbourhood of $1/2$ in $[0,1]$. We know that $s+s^{\prime}$ restricted to $(a,1/2]$ and to $[1/2,b)$ is in both cases an embedding. Which implies that $s+s^{\prime}(1/2)$ is going to be an internal point in $(s+s^{\prime})((a,b))$, which is then an open set. Thus $s+s^{\prime}$ is an homeomorphism also if restricted to $(a,b)$. Associativity follows easily.
In case $G$ is a directed graph we are done as considering $s^{\prime\prime}\sim s$ and $s^{\prime\prime\prime}\sim s^{\prime}$ would still give:
  • $s(1)=s^{\prime}(0)$;
  • $Im(s)\cap Im(s^{\prime})=s(1)$.
In case $G$ is an undirected graph, then $Im(s^{\prime\prime})\cap Im(s^{\prime\prime\prime})=s^{\prime\prime}(1)=s(1)$ holds in any case. However we may have $s^{\prime\prime}(1)\neq s^{\prime\prime\prime}(0)$. But, if that is the case, it is enough to change $s^{\prime\prime}(t)$ with $s^{\prime\prime}(1-t)$ and/or $s^{\prime\prime\prime}(t)$ with $s^{\prime\prime\prime}(1-t)$.
$\blacksquare$
Proof of 1.
To lighten the notation we use the following symbols:
  • the symbols $e_{d}$ and $e^{\prime-1}_{d}$ respectively mean that the edge $e$ is deleted and $e^{\prime}$ is inserted;
  • if instead we write $s_{g}$ and $s^{\prime-1}_{g}$ we imply that the segment $s$ is involved in a ghosting operation and the segment $s^{\prime}$ is the result of the splitting of an edge;
  • lastly, the shrinking induced by $(e,e^{\prime})$ is called $e_{w,w^{\prime}}$ with $w$ being the original weight function, and $w^{\prime}$ the weight function after the shrinking.
Note that all edits are defined on a single edge a part from ghostings and splittings, which in general are associated to a segment.
Suppose that $\gamma$ is a finite edit path. This means that $\gamma$ is the composition of a finite set of edits. We indicate such ordered composition with $\gamma=\prod_{i=0}^{N}e_{i}$ with $e_{i}$ edit operation. We would like to change the order of the edit operations without raising the cost and changing the extremes of the edit path. This is not always possible. However we can work it around in the useful cases using properties (P1)-(P4). In particular, we would like to know when we can commute a generic edit $e_{i}$ in the following situations:
  • $e_{d}\circ e_{i}$ and $e_{i}\circ e^{-1}_{d}$
  • $s_{g}\circ e_{i}$ and $e_{i}\circ e^{-1}_{g}$.
Moreover we want to reduce the edit path to max one shrinking/deletion/insertion for any edge of $G$ and $G^{\prime}$.
Inverse Operations
We point out the following fact which will be used in the proof. Suppose we have a couple of edits $e_{i}\circ e_{j}$ that can be replaced with the couple $e^{\prime}_{i}\circ e^{\prime}_{j}$. Now consider $e^{-1}_{j}\circ e^{-1}_{i}$:
$e_{d}$ and $e_{d}^{-1}$
Note that, when we delete or insert one edge $e$, we are modifying the graph only at the level of its open star $\mathcal{N}(e)$. Consider then the following scenarios:
  • $e_{d}\circ s_{g}$ so that the ghostings induced by $s$ result in the edge $e$, which is then deleted. This composition of edits can be safely replaced with the deletions of all the edges in $s$, conserving the length of the path by (P3). If the ghostings defined by $s$ do not involve $e$ then we can safely commute the operations.
  • $e_{d}\circ s_{g}^{-1}$, with $e$ being a new edge originated by the splitting induced by $s_{g}^{-1}$. This composition of edits can clearly be replaced with a shrinking of the original edge and then a suitably defined splitting (which avoids the edge $e$) without changing the path length - by (P3) and (P4).
  • $e_{d}\circ e_{w,w^{\prime}}$ can be replaced by $e_{d}$ potentially diminishing the length of the path, but surely not raising it by (P1).
  • $e_{w,w^{\prime}}\circ e_{d}^{-1}$ can be replaced with an insertion directly with weight $w(e)$, possibly shortening the path (P1).
  • $s^{\prime}_{g}\circ e_{d}^{-1}$follows from Equation 8.
  • $s^{\prime-1}_{g}\circ e_{d}^{-1}$ follows from Equation 8.
$s_{g}$ and $s_{g}^{-1}$
Like in the previous case, we only take into account transformations regarding edges in the open star of the ghostings/splittings.
  • $s_{g}\circ s^{\prime}_{g}$, with $s$ and $s^{\prime}$ being segments whose interiors have non-empty intersection, can be merged in a unique ghosting by (P2). Note that, in case $G$ and $G^{\prime}$ are directed graphs, then the directions of the two segments are compatible. Otherwise the intersection would be empty. We avoid repeating this same comment in the upcoming points.
  • $s_{g}\circ s^{\prime-1}_{g}$, with $s$ and $s^{\prime}$ being segments whose interiors have non-empty intersection, can commute provided we define carefully the splittings $s^{\prime-1}_{g}$ by (P2), - avoiding the splittings which lie in the intersection.
  • $s_{g}\circ e_{w,w^{\prime}}$ with $e\subset Im(s)$. This means that we are shrinking an edge before merging it with others via some ghostings. However, we can achieve the same result, without increasing the path length, by making the ghostings at first, and then shrinking the resulting edge - by (P1)-(P4).
  • $e_{w,w^{\prime}}\circ s_{g}^{-1}$ follows from Equation 8.
$e_{w,w^{\prime}}$
  • $v_{w^{\prime},w^{\prime\prime}}\circ e_{w,w^{\prime}}$ can be replaced by $e_{w,w^{\prime\prime}}$ which is either conserving or shortening the path (P1).
  • note that $v_{w,w^{\prime}}^{-1}=v_{w^{\prime},w}$.
Thanks to these properties we can take a given path $\gamma=\prod_{i=0,...,N}e_{i}$ and modify the edit operations in order to obtain the following situation:
  • the first operations are all in the form $e_{d}$; this can be achieved because $e_{d}\circ-$ can be always rearranged so that the deletions are always done first, potentially shortening the path. Of course there can be only one deletion for each edge of $G$;
  • then we have all the edits in the form $s_{g}$; since $s_{g}\circ-$ is exchangeable when we do not have $s_{g}$ pre-composed by some deletion, this is not a problem. Observe that all ghostings (and so all splittings) involving segments whose interiors have non empty intersection can be joined in a unique ghosting operation;
  • in the same way we can put last all the paths in the form $e_{d}^{-1}$ and before them the ones in the form $s_{g}^{-1}$;
  • in the middle we are left with the shrinking paths, which by construction are on edges which were segments in the original graphs. Since we can substitute $e_{w,w^{\prime}}\circ v_{w^{\prime},w^{\prime\prime}}$ with $e_{w,w^{\prime\prime}}$, we can obtain just one single shrinking for each edge and this means that the shrinkings are pairing uniquely segments in the original graphs.
Thus the edit path
$\overline{\gamma}=(\gamma_{d}^{G^{\prime}})^{-1}\circ(\gamma_{g}^{G^{\prime}})^{-1}\gamma_{s}^{G}\circ\gamma_{g}^{G}\circ\gamma_{d}^{G},$
with:
  • $\gamma_{d}^{G}=\prod e_{d}$
  • $\gamma_{g}^{G}=\prod s_{g}$
  • $\gamma_{s}^{G}=\prod e_{w,w^{\prime}}$
  • $(\gamma_{g}^{G^{\prime}})^{-1}=\prod s_{g}^{-1}$
  • $(\gamma_{d}^{G^{\prime}})^{-1}=\prod e_{d}^{-1}$
is such that $\gamma(T)=\overline{\gamma}(G)=G^{\prime}$ and $cost(\overline{\gamma})\leq cost(\gamma)$. The key point is that $\overline{\gamma}$ can be easily realized as a mapping whose couples are given by the segments paired by the shrinkings.
$\blacksquare$
Proof of 6.
@@
Let $D_{S}$ be its display poset and $(T_{S},f_{S}:V\rightarrow\mathbb{R})$ with $T_{S}=(V,E)$ its merge tree. By construction $V-\{r_{T_{S}}\}$ with $r_{T_{S}}$ being the root of the merge tree and $f_{S}(r_{T_{S}})=+\infty$ can be sent into $D_{S}$: every vertex $v$ in $V-\{r_{T_{S}}\}$ is identified by a critical value $t_{i}=f_{S}(v)$ and an element $v\in S(t_{i})$. Thus we have $V-\{r_{T_{S}}\}\hookrightarrow D_{S}$. Moreover, this embedding is an embedding of posets if we consider $T_{S}$ with the relationship generated by $child<father$.
@@
This implies that for any edge $e$ connecting $(t,s)\rightarrow(t_{j},p)$ in the DAG representing $D_{S}$, we have that if $(t,s)\neq(t_{i},S(t_{i}\leq t_{j})^{-1}(p))$, then we also have an edge $(t,s)\rightarrow(t_{i},S(t_{i}\leq t_{j})^{-1}(p))$. But this implies that we can remove $e$ from the DAG without changing its transitive closure. That is, $e$ does not appear in $\mathcal{G}(D_{S})$. The same for any edge $(t_{j},p)\rightarrow(t,s)$ with $(t,s)\neq(t_{k},S(t_{i}\leq t_{k})(p))$. Thus $(t_{j},p)$ induces an order $2$ vertex in $\mathcal{G}(D_{S})$.
To conclude the proof we just need to prove that $\mathcal{G}(D_{S})$ is a tree. In fact, if this is the case, then we have a poset map $V-\{r_{T}\}\hookrightarrow V_{\mathcal{G}(D_{S})}$ so that all the vertices not in the image of this map are of order $2$. I.e. the map is an isomorphism of posets onto the image, which is the poset obtained by removing all order $2$ vertices from $\mathcal{G}(D_{S})$.
@@
Any order $2$ vertex which is not ghosted is paired with another order $2$ vertex. Ghosting both of them does not increase the cost of the mapping.
@@
Proof of 7.
@@
Condition (M2) coincide with condition (A2). Condition (M3) is clearly satisfied because of the antichain condition (A1). Consider a vertex $v\in E_{T}$. The only case in which $v$ is not edited is when $v<x$ with $x\in v_{T}\cap\pi_{T}(M^{*})$. However, in this case, $v$ does not appear in $\widetilde{T}$, and thus (M1) is satisfied. Moreover, all and only order $2$ vertices, after the deletions, are ghosted, and (M4) follows .
@@
We note that $father>son$ induces a partial order relationship also on the pairs given by coupled points in $M$: $(x,y)>(v,w)$ if $x>v$ and $y>w$. In fact, by property (M3), $x>v$ if and only if $y>w$. So we can select $(x_{i},y_{i})$, the maxima with respect to this partial order relationship. Thus, we obtain $(x_{0},y_{0})$,…,$(x_{n},y_{n})$ which form an antichain (both in $V_{T}$ and $V_{T^{\prime}}$).
Clearly $M^{*}=\{(x_{0},y_{0})$,…,$(x_{n},y_{n})\}\in\mathcal{C}^{*}(T,T^{\prime})$. Now we build $\alpha(M^{*})$ and compare the cost of its edits with the ones in $M$. Let $\bar{x}=LCA(x_{i},x_{j})$. Since $\bar{x}>x_{i},x_{j}$, it is not coupled in $M$. Since $x_{i}$ and $x_{j}$ are coupled, $\bar{x}$ cannot be ghosted, so it is deleted in $M$. Any point $x$ above $\bar{x}$ is deleted for the same reasons. So the edits above $\bar{x}$ are shared between $\alpha(M^{*})$ and $M$.
@@
by properties (P1)-(P4). Since $\alpha(M^{*})\in M_{2}(\widetilde{T},\widetilde{T}^{\prime})$ by 7, we have:
@@
Proof of 8.
Having fixed a leaf $l$, the constraint in Equation (4) allows for at most one path $\zeta_{v}^{v_{i+1}}\subset\zeta_{l}$ to be kept after the edits induced by all the variables equal to $1$. Moreover if $(v^{\prime\prime},i)\in\Phi(v)\cap\Phi(v^{\prime})$, then $v=v^{\prime\prime}_{i}=v^{\prime}$. Thus, variables are added at most one time in Equation 4 and Equation 5. Which means that for any $a\in V_{T_{x}}$, we are forcing that the vertex $a$ can be an internal vertex or lower extreme of at most one path $\zeta_{v}^{v_{i+1}}$ such that $\delta_{i,j}^{v,w}=1$. In other words if two “kept” segments $\zeta_{v}^{v_{i+1}}$ and $\zeta_{v^{\prime}}^{v^{\prime}_{i^{\prime}+1}}$ (i.e. with $\delta_{i,j}^{v,w}=\delta_{i^{\prime},j}^{v^{\prime},w}=1$) intersect each other, it means that they just share the upper extreme $v_{i+1}=v^{\prime}_{i^{\prime}+1}$. These facts together imply that (if the constraints are satisfied) the edits induced on $T_{x}$ by $\delta^{v,w}_{i,j}=1$ and $\delta^{v^{\prime},w^{\prime}}_{i^{\prime},j^{\prime}}=1$ are always compatible: if $v^{\prime\prime}\in V_{sub_{T}(v_{i})}$ then it is not touched by (the edits induced by) $\delta^{v^{\prime},w^{\prime}}_{i^{\prime},j^{\prime}}=1$ (and the same exchanging the role of $v^{\prime}$ and $v$), if $v^{\prime\prime}$ is equal or above $v_{i+1}$ and/or $v^{\prime}_{i^{\prime}+1}$, then it is deleted in any case. Lastly, by noticing that if $\delta^{v,w}_{i,j}\in\Phi(v^{\prime})$ then $\delta^{v,w^{\prime}}_{i,j^{\prime}}\in\Phi(v^{\prime})$ for all other possible $w^{\prime}$ and $j^{\prime}$, we see that every segment $\zeta_{v}^{v_{i}}$ is paired with at most one segment $\zeta_{w}^{w_{j}}$, and viceversa.
As a consequence, for any vertex $v^{\prime}$ in any of the tree structures, at most one point on the segment $\zeta_{v^{\prime}}$ is coupled in $M^{*}$, guaranteeing the antichain condition. Moreover, any point of $T_{x}$ which is in $\pi_{T_{x}}(M^{*})$ is assigned to one and only one point of $T^{\prime}_{y}$ and viceversa. The edits induced by $\delta=1$ clearly satisfy properties (M2)-(M4). Passing to $\widetilde{T}_{x}$ and $\widetilde{T}_{y}$, also (M1) is satisfied.

### v4

@@ B Proofs
$W_{1}([D],[D^{\prime}])=W_{1}(\sum_{x\in[D]}\{x\},\sum_{y\in[D^{\prime}]}\{y\})=\sup_{k\in\text{Lip}(\mathbb{R}^{2}_{\Delta})}(\sum_{x\in[D]}k(x)-\sum_{y\in[D^{\prime}]}k(y)).$
@@
  • replacing $[D^{\prime}]$ with the empty diagram and noting:
    $\sup_{k\in\text{Lip}(\mathbb{R}^{2}_{\Delta})}(\sum_{x\in[D]}k(x))=\sum_{x\in[D]}\min_{y\in\Delta}\parallel x-y\parallel_{1};$
  • observing that when we have $W_{1}([D+D^{\prime\prime}],[D^{\prime}+D^{\prime\prime}])$ the contributions by the elements in $[D^{\prime\prime}]$ when computing:
    $\sum_{x\in[D+D^{\prime\prime}]}k(x)-\sum_{y\in[D^{\prime}+D^{\prime\prime}]}k(y)$
@@
Proof of 5.
@@
Let $D_{S}$ be its display poset and $(T_{S},f_{S}:V\rightarrow\mathbb{R})$ with $T_{S}=(V,E)$ its merge tree. By construction $V-\{r_{T_{S}}\}$ with $r_{T_{S}}$ being the root of the merge tree and $f_{S}(r_{T_{S}})=+\infty$ can be sent into $D_{S}$: every vertex $v$ in $V-\{r_{T_{S}}\}$ is identified by a critical value $t_{i}=f_{S}(v)$ and an element $v\in S(t_{i})$. Thus we have $V-\{r_{T_{S}}\}\hookrightarrow D_{S}$. Moreover, this embedding is an embedding of posets if we consider $T_{S}$ with the relation generated by $child<parent$.
@@
This implies that for any edge $e$ connecting $(t,s)\rightarrow(t_{j},p)$ in the DAG representing $D_{S}$, we have that if $(t,s)\neq(t_{i},S(t_{i}\leq t_{j})^{-1}(p))$, then we also have an edge $(t,s)\rightarrow(t_{i},S(t_{i}\leq t_{j})^{-1}(p))$. But this implies that we can remove $e$ from the DAG without changing its transitive closure. That is, $e$ does not appear in $\mathcal{G}(D_{S})$. The same for any edge $(t_{j},p)\rightarrow(t,s)$ with $(t,s)\neq(t_{k},S(t_{i}\leq t_{k})(p))$. Thus $(t_{j},p)$ induces a degree $2$ vertex in $\mathcal{G}(D_{S})$.
To conclude the proof we just need to prove that $\mathcal{G}(D_{S})$ is a tree. In fact, if this is the case, then we have a poset map $V-\{r_{T}\}\hookrightarrow V_{\mathcal{G}(D_{S})}$ so that all the vertices not in the image of this map are of degree $2$. I.e. the map is an isomorphism of posets onto the image, which is the poset obtained by removing all degree $2$ vertices from $\mathcal{G}(D_{S})$.
@@
Any degree $2$ vertex which is not ghosted is paired with another degree $2$ vertex. Ghosting both of them does not increase the cost of the mapping.
@@
Proof of 2.
To lighten the notation we use the following symbols:
  • the edit induced by $(v,\mathfrak{D})$ is called $v_{d}$ and $v^{-1}_{d}$ stands for $(\mathfrak{D},v)$.
  • the edit induced by $(v,\mathfrak{G})$ is called $v_{g}$ and $v^{-1}_{g}$ stands for $(\mathfrak{G},v)$.
  • the edit induced by $(v,v^{\prime})$ is called $v_{\varphi,\varphi^{\prime}}$ with $\varphi$ being the original weight function, and $\varphi^{\prime}$ the weight function after the shrinking.
We know that the set of finite edit paths between two dendrograms is nonempty.
Suppose that $\gamma$ is a finite edit path. This means that $\gamma$ is the composition of a finite set of edits. We indicate such ordered composition with $\gamma=\prod_{i=0}^{N}e_{i}$ with $e_{i}$ edit operation. We would like to change the order of the edit operations without raising the cost and changing the endpoints of the edit path. This is not always possible. However we can work it around in the useful cases using properties (P1)-(P4). In particular, we would like to know when we can commute a generic edit $e_{i}$ in the following situations:
  • $v_{d}\circ e_{i}$ and $e_{i}\circ v^{-1}_{d}$
  • $v_{g}\circ e_{i}$ and $e_{i}\circ v^{-1}_{g}$.
Moreover we want to reduce the edit path to max one edit for any vertex of $T$ and $T^{\prime}$.
Inverse Operations
We point out the following fact, which will be used often in the proof. Suppose we have a pair of edits $e_{i}\circ e_{j}$ that can be replaced with the pair $e^{\prime}_{i}\circ e^{\prime}_{j}$. Now consider $e^{-1}_{j}\circ e^{-1}_{i}$:
$v_{d}$ and $v_{d}^{-1}$
When we delete or insert one vertex, we are modifying the tree structure at the level of its parent and its children. Therefore, we are only taking into account operations involving the parent of the considered vertex, on the vertex himself or on the children of the deleted/inserted vertex.
  • $v_{d}\circ v^{\prime}_{g}$: if $v$ is a child of $v^{\prime}$, can be safely replaced with $v_{d}\circ v^{\prime}_{d}$. Instead of ghosting the parent and then deleting the whole edge, we can delete both edges one by one; conserving the length of the path (P3). If, instead, $v$ is parent of $v^{\prime}$, we can safely commute the operations.
  • $v_{d}\circ v_{g}^{-1}$: the two edits can be replaced with $v^{\prime}_{\varphi,\varphi^{\prime}}$ with $v^{\prime}$ parent of $v$ (after the insertion) and $\varphi^{\prime}$ properly defined not to raise the cost of the path. In fact we are inserting $v$ on an edge and then deleting it. This can obviously be achieved by shrinking the original edge (without changing the path length - (P4)).
  • $v_{d}\circ v^{\prime-1}_{g}$: if $v^{\prime}$ is parent of $v$ after the splitting, we can again replace the two edits with shrinking: instead of inserting a point in an edge, and deleting then the edge below, we can directly shrink the original edge (P4). If, instead, $v^{\prime}$ is inserted below $v$, we have the same situation, but seen from the point of view of the vertex which is the child of $v^{\prime}$ after the splitting.
  • $v_{d}\circ v_{\varphi,\varphi^{\prime}}$: the two edits can be replaced by $v_{d}$ potentially diminishing the length of the path, but surely not raising it (P1).
  • $v^{\prime}_{g}\circ v_{d}^{-1}$: if $v^{\prime}$ is the parent of $v$, this edit can be replaced with just $v^{\prime}_{\varphi,\varphi^{\prime}}$ with appropriate weights: we are inserting an edge under a vertex which (in this case) becomes of degree two and is ghosted. We can directly modify the edge without changing the length of the path (P4). If, instead, $v^{\prime}$ becomes a child of $v$ after the insertion, we can simply shrink $v^{\prime}$ to obtain the same result without raising the cost (P4).
  • $v^{\prime-1}_{g}\circ v_{d}^{-1}$: if $v^{\prime}$ is on the edge inserted with $v_{d}^{-1}$, we cannot commute these two edits, but we can replace them with two insertions: instead of inserting an edge and then splitting it, we can directly insert two smaller edges; without changing the cost of the path (P3). In all other situations, we can commute the two operations.
  • $v_{\varphi,\varphi^{\prime}}\circ v_{d}^{-1}$: the two edits can be replaced with an insertion directly with weight $\varphi^{\prime}$, possibly shortening the path (P1).
  • $v^{\prime-1}_{d}\circ v_{d}$: suppose that the parent of $v^{\prime}$, after its insertion, is the parent of $v$ before we apply both edits. If the children of $v^{\prime}$ are different from the children of $v$, this operations do not commute. If the children are the same, they can be changed with a shrinking of $v$, reducing the length of the path by at most $cost(v^{\prime-1}_{d})+cost(v_{d})$ (P1).
$v_{g}$ and $v_{g}^{-1}$
Like in the previous case, we only take into account transformations concerning the parent and the child of the added/ghosted degree two vertex.
  • $v_{g}\circ v^{\prime}_{g}$: always commute (P2).
  • $v_{g}\circ v^{\prime-1}_{g}$: always commute, provided that we define carefully the splitting $v^{\prime-1}_{g}$ (P2).
  • $v_{g}\circ v_{\varphi,\varphi^{\prime}}$: with such operations, we are shrinking a vertex before ghosting it. However, we can achieve the same result, without increasing the path length, by ghosting the vertex at first, and then shrinking its child (P1)-(P4).
  • $v^{\prime}_{\varphi,\varphi^{\prime}}\circ v_{g}^{-1}$: if either $v^{\prime}=v$, or $v$ is the parent of $v^{\prime}$, we can replace those edits with an appropriate shrinking of the vertex which is the child of $v$ after the splitting and an appropriate insertion of $v^{\prime}$, without changing the length of the path (P3)-(P4).
  • $v_{g}\circ v^{\prime}_{d}$: if $v$ is the parent of $v^{\prime}$, such edits do not commute and can’t be replaced by a similar operation which inverts ghosting and deletion.
$v_{w,w^{\prime}}$
  • $v_{\varphi^{\prime},\varphi^{\prime\prime}}\circ v_{\varphi,\varphi^{\prime}}$: those edit can be replaced by $v_{\varphi,\varphi^{\prime\prime}}$ which is either conserving or shortening the path (P1).
  • $v_{\varphi,\varphi^{\prime}}^{-1}$: we can proceed as in the previous point, since: $v_{\varphi,\varphi^{\prime}}^{-1}=v_{\varphi^{\prime},\varphi}$.
Thanks to these properties, we can take a given path $\gamma=\prod_{i=0,...,N}e_{i}$ and modify the edit operations in order to obtain the following situation:
  • the first operations are all in the form $v_{d}$; this can be achieved because $v_{d}\circ-$ can be always rearranged, potentially by changing the path (as shown before) and shortening it. Of course, there can be only one deletion for each vertex of $T$;
  • then we have all the edits in the form $v_{g}$; since $v_{g}\circ-$ is exchangeable any time but when we have $v_{g}\circ v^{\prime}_{d}$, this is not a problem. Observe that all degree two vertices which were not deleted can be ghosted (at most one time);
  • in the same way we can put last all the paths in the form $v_{d}^{-1}$ and before them $v_{g}^{-1}$. All the new vertices appearing with the insertion of edges and the splitting of edges with degree two vertices are all nodes which remain in $T^{\prime}$ and which are not further edited;
  • in the middle we are left with the shrinking paths. Since we can substitute $v_{\varphi,\varphi^{\prime}}\circ v_{\varphi^{\prime},\varphi^{\prime\prime}}$ with $v_{\varphi,\varphi^{\prime\prime}}$, we can obtain at most one transformation on each vertex.
Thus, we can find at least one edit path of the form:
$\overline{\gamma}=(\gamma_{d}^{T^{\prime}})^{-1}\circ(\gamma_{g}^{T^{\prime}})^{-1}\gamma_{s}^{T}\circ\gamma_{g}^{T}\circ\gamma_{d}^{T}$
with:
  • $\gamma_{d}^{T}=\prod v_{d}$;
  • $\gamma_{g}^{T}=\prod v_{g}$;
  • $\gamma_{s}^{T}=\prod v_{\varphi,\varphi^{\prime}}$;
  • $(\gamma_{g}^{T^{\prime}})^{-1}=\prod v_{g}^{-1}$;
  • $(\gamma_{d}^{T^{\prime}})^{-1}=\prod v_{d}^{-1}$.
As for the paths $\gamma_{M}$ described in Section 4.2, all the permutations of the edits inside each $\gamma_{d}^{T},\gamma_{g}^{T},\ldots,$ does not change its endpoints and its cost. Thus, $\overline{\gamma}$ is such that $\gamma(T)=\overline{\gamma}(T)=T^{\prime}$ (up to isomorphism) and $cost(\overline{\gamma})\leq cost(\gamma)$.
The key point is that $\overline{\gamma}$ can be easily realized as a mapping made by the following elements:
  • $(v,\mathfrak{D})$, $\forall v_{d}\in\gamma_{d}^{T}$;
  • $(v,\mathfrak{G})$, $\forall v_{g}\in\gamma^{T}_{g}$;
  • $(v,v^{\prime})$, $\forall v_{\varphi,\varphi^{\prime}}\in\gamma_{s}^{T}$, where $v^{\prime}$ is the edge in $T^{\prime}$ (more precisely, in $\gamma_{g}^{T^{\prime}}\circ\gamma_{d}^{T^{\prime}}(T^{\prime})$) associated to the shrinking $v_{\varphi,\varphi^{\prime}}$;
  • $(\mathfrak{G},v)$, $\forall v^{-1}_{g}\in(\gamma_{g}^{T^{\prime}})^{-1}$;
  • $(\mathfrak{D},v)$, $\forall v^{-1}_{d}\in(\gamma_{d}^{T^{\prime}})^{-1}$.
$\blacksquare$
Proof of 6.
@@
Condition (M2) coincide with condition (A2). Condition (M3) is clearly satisfied because of the antichain condition (A1). Consider a vertex $v\in E_{T}$. The only case in which $v$ is not edited is when $v<x$ with $x\in v_{T}\cap\pi_{T}(M^{*})$. However, in this case, $v$ does not appear in $\widetilde{T}$, and thus (M1) is satisfied. Moreover, after the deletions, all degree $2$ vertices are ghosted, and (M4) follows.
@@
We note that $parent>child$ induces a partial order relation also on the pairs given by paired points in $M$: $(x,y)>(v,w)$ if $x>v$ and $y>w$. In fact, by property (M3), $x>v$ if and only if $y>w$. So we can select $(x_{i},y_{i})$, the maxima with respect to this partial order relation . Thus, we obtain $(x_{0},y_{0})$,…,$(x_{n},y_{n})$ which form an antichain (both in $V_{T}$ and $V_{T^{\prime}}$).
Clearly $M^{*}=\{(x_{0},y_{0})$,…,$(x_{n},y_{n})\}\in\mathcal{C}^{*}(T,T^{\prime})$. Now we build $\alpha(M^{*})$ and compare the cost of its edits with the ones in $M$. Let $\bar{x}=LCA(x_{i},x_{j})$. Since $\bar{x}>x_{i},x_{j}$, it is not paired in $M$. Since $x_{i}$ and $x_{j}$ are paired, $\bar{x}$ cannot be ghosted, so it is deleted in $M$. Any point $x$ above $\bar{x}$ is deleted for the same reasons. So the edits above $\bar{x}$ are shared between $\alpha(M^{*})$ and $M$.
@@
by properties (P1)-(P4). Since $\alpha(M^{*})\in M_{2}(\widetilde{T},\widetilde{T}^{\prime})$ by 6, we have:
@@
Proof of 7.
Having fixed a leaf $l$, the constraint in Equation (5) allows for at most one path $\zeta_{v}^{v_{i+1}}\subset\zeta_{l}$ to be kept after the edits induced by all the variables equal to $1$. Moreover if $(v^{\prime\prime},i)\in\Phi(v)\cap\Phi(v^{\prime})$, then $v=v^{\prime\prime}_{i}=v^{\prime}$. Thus, variables are added at most one time in Equation 5 and Equation 6. Which means that for any $a\in V_{T_{x}}$, we are forcing that the vertex $a$ can be an internal vertex or lower extreme of at most one path $\zeta_{v}^{v_{i+1}}$ such that $\delta_{i,j}^{v,w}=1$. In other words if two “kept” paths $\zeta_{v}^{v_{i+1}}$ and $\zeta_{v^{\prime}}^{v^{\prime}_{i^{\prime}+1}}$ (i.e. with $\delta_{i,j}^{v,w}=\delta_{i^{\prime},j}^{v^{\prime},w}=1$) intersect each other, it means that they just share the upper extreme $v_{i+1}=v^{\prime}_{i^{\prime}+1}$. These facts together imply that (if the constraints are satisfied) the edits induced on $T_{x}$ by $\delta^{v,w}_{i,j}=1$ and $\delta^{v^{\prime},w^{\prime}}_{i^{\prime},j^{\prime}}=1$ are always compatible: if $v^{\prime\prime}\in V_{sub_{T}(v_{i})}$ then it is not touched by (the edits induced by) $\delta^{v^{\prime},w^{\prime}}_{i^{\prime},j^{\prime}}=1$ (and the same exchanging the role of $v^{\prime}$ and $v$), if $v^{\prime\prime}$ is equal or above $v_{i+1}$ and/or $v^{\prime}_{i^{\prime}+1}$, then it is deleted in any case. Lastly, by noticing that if $\delta^{v,w}_{i,j}\in\Phi(v^{\prime})$ then $\delta^{v,w^{\prime}}_{i,j^{\prime}}\in\Phi(v^{\prime})$ for all other possible $w^{\prime}$ and $j^{\prime}$, we see that every path $\zeta_{v}^{v_{i}}$ is paired with at most one path $\zeta_{w}^{w_{j}}$, and viceversa.
As a consequence, for any vertex $v^{\prime}$ in any of the tree structures, at most one point on the path $\zeta_{v^{\prime}}$ is paired in $M^{*}$, guaranteeing the antichain condition. Moreover, any point of $T_{x}$ which is in $\pi_{T_{x}}(M^{*})$ is assigned to one and only one point of $T^{\prime}_{y}$ and viceversa. The edits induced by $\delta=1$ clearly satisfy properties (M2)-(M4). Passing to $\widetilde{T}_{x}$ and $\widetilde{T}_{y}$, also (M1) is satisfied.